Cost, revenue, and profit are in dollars and x is the number of units. The average cost of a product changes at the rate C(xCost, revenue, and profit are in dollars and x is the number of units. If the daily marginal cost for a product is MC = 8x +please do all three parts:)thanks a ton
Answer #1

a) 7 C) = Pēcas de Ĉ (x) = -8 + ata 8 a is a constanta to be determined - 2+1 2+1 - č. (x) = 8 + z ta Now we know ē (8) = 12ī (16)= = 8 + +10 {+ 2+10 = { +12= 229 [ (16)= $ 12.5! Ĉ (16) = $12 and 50 cents! Mc = 8x+120 Total cost fonction = (18x+120)